# SuPreME: A Supervised Pre-training Framework for Multimodal ECG Representation Learning

### v1

@@ Abstract
Cardiovascular diseases are a leading cause of death and disability worldwide. Electrocardiogram (ECG) recordings are critical for diagnosing and monitoring cardiac health, but obtaining large-scale annotated ECG datasets is labor-intensive and time-consuming. Recent ECG Self-Supervised Learning (eSSL) methods mitigate this by learning features without extensive labels but fail to capture fine-grained clinical semantics and require extensive task-specific fine-tuning. To address these challenges, we propose SuPreME, a Supervised Pre-training framework for Multimodal ECG representation learning. SuPreME applies Large Language Models (LLMs) to extract structured clinical entities from free-text ECG reports, filter out noise and irrelevant content, enhance clinical representation learning, and build a high-quality, fine-grained labeled dataset. By using text-based cardiac queries instead of traditional categorical labels, SuPreME enables zero-shot classification of unseen diseases without additional fine-tuning. We evaluate SuPreME on six downstream datasets covering 127 cardiac conditions, achieving superior zero-shot AUC performance over state-of-the-art eSSL and multimodal methods by over 1.96% [^1] All code and data will be released upon acceptance.. Results demonstrate the effectiveness of SuPreME in leveraging structured, clinically relevant knowledge for high-quality ECG representations.
@@ 1 Introduction
To reduce dependence on annotations, recent advancements in ECG self-supervised learning (eSSL) have enabled the extraction of representative features from large-scale unannotated ECGs using contrastive or generative tasks (Eldele et al., 2021; Kiyasseh et al., 2021; Na et al., 2024). Despite their promise, these methods often rely on strong signal-level augmentations that may distort the semantic integrity of the signal and require complex pretext task designs (Liu et al., 2024; Kiyasseh et al., 2021). Multimodal learning approaches (Liu et al., 2024; Li et al., 2024) have also been proposed to learn ECG representations by leveraging free-text ECG reports. However, these methods face challenges due to noise in textual data and the complexities of language grammar, which can hinder learning efficiency (Wu et al., 2023).
Figure 1: Overview of SuPreME framework, including (a) ECG clinical entity extraction with LLMs, (b) pre-training framework for SuPreME, and (c) zero-shot prompted evaluation.
To address these limitations and develop a scalable, simple, and effective ECG pre-training framework, we propose SuPreME, a Supervised Pre-training framework for Multimodal ECG representation learning, as illustrated in Figure 1. Our contributions are threefold: (a) We introduce an automated pipeline for extracting high-quality clinical entities from raw ECG reports via LLMs and professional medical databases, yielding a fine-grained ECG dataset with 295 standardized medical terminologies. This approach requires no manual labeling, ensuring both scalability and consistency while capturing richer semantics than coarse-grained or free-text labels. (b) Building on these entities, we propose SuPreME, which avoids data augmentation and complex pretext tasks by directly aligning ECG signals with structured entity labels. This approach surpasses prior multimodal methods (e.g., MERL) that rely on raw free-text, underscoring the benefits of explicit entity-level supervision. (c) We train SuPreME on 771,500 ECG signals paired with extracted entities from the MIMIC-IV-ECG dataset (Gow et al., 2023). On six downstream datasets, it significantly outperforms state-of-the-art multimodal methods (Liu et al., 2024) in zero-shot scenario, even those relying on human-assisted prompt engineering. SuPreME also surpasses fully fine-tuned eSSLs using only 20% of the pre-training data in a purely zero-shot setting, underscoring its simplicity, effectiveness, and high data efficiency.
@@ 2 Related Work
Biomedical Multimodal Learning. Multimodal learning has advanced significantly in biomedical applications, especially in vision-language pre-training (VLP) frameworks for radiology (Liu et al., 2023b, a; Wan et al., 2024; Zhang et al., 2023b; Wu et al., 2023), which align radiology images with structured knowledge from reports to reduce noise and improve robustness. However, multimodal learning for ECG remains underexplored. Methods like MERL (Liu et al., 2024) and ECG-LM (Yang et al., ) integrate ECG signals and raw text reports but struggle with noise and inconsistencies in unstructured reports. Others, such as KED (Tian et al., 2024), use structured labels and contrastive learning strategies but face challenges from label noise and LLM-generated knowledge hallucinations. Our approach addresses these issues by structuring reports into meaningful entities, reducing noise, and aligning them with ECG signals without reliance on LLM-augmented content, minimizing hallucination risks while enabling efficient representation learning and downstream flexibility.
@@ 3 Methodology
SuPreME extracts structured clinical entities from ECG reports via LLMs (Section 3.1) and embeds both ECG signals and text-based cardiac queries into a shared space. A Cardiac Fusion Network (CFN, Section 3.2) then aligns these embeddings, enabling zero-shot classification of unseen cardiac conditions without fine-tuning (Section 3.3), yielding scalable, clinically meaningful representations.
3.1 ECG Clinical Entity Extraction
Enriching LLM with Domain-specific Knowledge. ECG reports, generated by 12-lead devices (see Appendix A.1), provide diverse insights into cardiac activity. To harness this diversity and enhance model performance, We utilize domain-specific knowledge, including ECG terminologies and abbreviations (e.g., "sinus rhythm", "LVH" for Left Ventricular Hypertrophy), compiled systematically with GPT-4o from clinician-validated databases like UMLS [^2] https://www.nlm.nih.gov/research/umls and SNOMED CT [^3] https://www.snomed.org/ (Bodenreider, 2004; Donnelly et al., 2006). SCP codes from datasets like PTB-XL further enrich this knowledge base.
Figure 3: Design of ECG report entity extraction with (a) knowledge-enhanced prompt engineering, and (b) candidate entity deduplication and mapping.
Knowledge-Guided ECG Entity Extraction. Before performing Named Entity Recognition (NER) with Llama3.1-70B-Instruct [^4] https://huggingface.co/meta-llama., we first provide domain-specific knowledge, including key concepts, terminologies, and diagnostic structures, ensuring the model internalizes relevant medical context. Few-shot examples from manually annotated clinical reports further refine its understanding. Once primed, the model extracts entities (e.g., waveforms, diagnoses) from free-text ECG reports, classifying them at a “Global” level into ‘Normal’, ‘Abnormal’, or ‘Uncertain’ (in Figure 3(a)) to filter unreliable diagnostics. Similar entities within a report are merged for consistency, laying the groundwork for cross-report deduplication.
Entity Deduplication with Professional Medical Database. Despite structured extraction and filtering, variations in physician writing styles and device-specific formats introduce redundancies and inconsistencies in the extracted entities. To resolve this, we compile a standardized ECG terminology library using LLMs to identify terms from professional medical databases like UMLS and SNOMED CT. We then use Medical Contrastive Pre-trained Transformers [^5] https://huggingface.co/ncbi/MedCPT-Query-Encoder. (MedCPT), a pre-trained medical BERT model, to align extracted entities with standardized terms (in Figure 3(b)). Using the PubMedBERT-initialized query encoder (QEnc) (Gu et al., 2021), cosine similarity between embeddings maps each entity to its closest standardized term. This deduplication and mapping process reduces 3,138 initial entities to 295 standardized terms, as shown in Figure 4, significantly enhancing consistency and ensuring alignment with professional medical terminology.
Figure 4: Frequency distribution of standardized ECG entities after deduplication and mapping.
@@ 3.2 Multimodal ECG Supervised Learning
ECG Embedding with Vision Transformer. The Vision Transformer (ViT) (Dosovitskiy et al., 2020), designed for 2D image processing, reshapes images into sequences of flattened patches for Transformer-based analysis. Similarly, ECG signals exhibit temporal and structural patterns analogous to the spatial relationships in images. We then adapt its architecture by dividing ECG time series into fixed-size patches, as shown in Figure 5.
Figure 5: ECG 1D ViT encoder in the SuPreME framework, with both lead-wise and position embedding.
@@
An ECG signal sequence is represented as a 2D matrix $\mathbf{x}\in\mathbb{R}^{L\times T}$, where $L$ is the number of leads and $T$ is the time series length. The time series is segmented into $N=T/P$ fixed-length patches, each containing $P$ time steps with $L$ leads, resulting in patches $\mathbf{x}_{p}\in\mathbb{R}^{L\times P}$. Each patch is flattened into a vector $\mathbf{x}_{p}\in\mathbb{R}^{L\cdot P}$, and concatenated to form a sequence $\mathbf{X}_{p}\in\mathbb{R}^{N\times(L\cdot P)}$. A learnable linear projection $\mathbf{W}_{p}\in\mathbb{R}^{(L\cdot P)\times D}$ maps each patch to the $D$-dimensional latent space:
$\mathbf{Z}_{p}=\mathbf{X}_{p}\cdot\mathbf{W}_{p},\mathbf{Z}_{p}\in\mathbb{R}^{N\times D}$ (1)
Similar to the $[class]$ token in BERT, we add a unique learnable embedding $\mathbf{e}_{i}\in\mathbb{R}^{D}$ for each lead $i$ and append it to the corresponding patch embedding $\mathbf{Z}_{p}$ to capture lead-specific features in the Transformer encoder. Positional embeddings $\mathbf{p}_{j}$ are added to preserve the order of patches in the time series. We use standard 1D positional embeddings $\mathbf{p}_{j}$ since it is sufficient to retain the sequential order in time series data. After incorporating $\mathbf{e}i$ and $\mathbf{p}j$, the final patch embedding $\mathbf{Z}{\text{final}}^{(i,j)}$ for each patch $j$ of lead $i$ is obtained. All embeddings $\mathbf{Z}{\text{final}}^{(i,j)}$ are concatenated along the lead dimension to form the sequence $\mathbf{Z}\in\mathbb{R}^{[B,L\times N,D]}$, where $B$ is the batch size.
$\begin{array}[]{c}\mathbf{Z}_{\text{final}}^{(i,j)}=\mathbf{Z}_{p}^{(i,j)}+\mathbf{e}_{i}+\mathbf{p}_{j}\\[5.0pt] \mathbf{Z}=\text{Concat}(\mathbf{Z}_{\text{final}}^{(1,1)},\dots,\mathbf{Z}_{\text{final}}^{(1,N)},\dots,\mathbf{Z}_{\text{final}}^{(L,N)})\end{array}$ (2)
The concatenated sequence $\mathbf{Z}$ is then processed through multiple Transformer blocks to extract high-level ECG features. Each block consists of self-attention and feed-forward layers with residual connections. To improve generalization, we incorporate stochastic depth dropout in the residual pathways. The detailed structure and mathematical formulation of the Transformer blocks are provided in Appendix A.2.
Cardiac Query Embedding with MedCPT. Rather than using traditional category labels, our framework adopts a flexible and semantically meaningful approach based on cardiac text queries, defined by the Standard Communication Protocol (SCP) system for ECG diagnoses. Descriptive queries are constructed using interpretations from medical databases like UMLS and SNOMED CT, ensuring consistency with the target representations in entity deduplication. Afterwards, MedCPT (Jin et al., 2023), a pre-trained medical BERT model, is employed to embed text queries into a shared latent space with ECG signal embeddings, using its query encoder (QEnc) initialized with PubMedBERT (Gu et al., 2021) in the retriever stage.
@@
$\mathbf{E}(q)=\mathbf{QEnc}(q)=\mathbf{Trm}([\text{CLS}]\ q\ [\text{SEP}])$ (3)
To ensure alignment with ECG signal embeddings for multimodal fusion, we apply a learnable projection layer that maps text query embeddings into the same latent space as the ECG features, as detailed in Appendix A.2. This enables effective cross-modal representation learning, facilitating the integration of structured clinical knowledge with ECG signals.
Alignment by Cardiac Fusion Network. The Cardiac Fusion Network (CFN) in SuPreME consists of multi-layer Transformer decoders, which align ECG signals with text-based cardiac queries by treating textual features as input queries and signal features as memory, as illustrated in Figure 6.
Figure 6: Architecture of the Cardiac Fusion Network (CFN) in the SuPreME.
Given a batch of signal features $\mathbf{F}_{\text{signal}}\in\mathbb{R}^{B\times N\times D}$ and text features $\mathbf{F}_{\text{text}}\in\mathbb{R}^{M\times D}$, where $B$ is the batch size, $N$ is the number of image blocks, $M$ is the number of text queries, and $D$ is the embedding dimension, the CFN learns a joint representation that captures cross-modal dependencies. The Transformer decoder attends to the ECG features while grounding the interpretation in medical queries. The fused output is processed by an MLP head for classification.
Further architectural details, including initialization, normalization, and implementation specifics, are described in Appendix A.2.

### v2

@@ Abstract
Cardiovascular diseases are a leading cause of death and disability worldwide. Electrocardiogram (ECG) is critical for diagnosing and monitoring cardiac health, but obtaining large-scale annotated ECG datasets is labor-intensive and time-consuming. Recent ECG Self-Supervised Learning (eSSL) methods mitigate this by learning features without extensive labels but fail to capture fine-grained clinical semantics and require extensive task-specific fine-tuning. To address these challenges, we propose SuPreME, a Supervised Pre-training framework for Multimodal ECG representation learning. SuPreME is pre-trained using structured diagnostic labels derived from ECG report entities through a one-time offline extraction with Large Language Models (LLMs), which help denoise, standardize cardiac concepts, and improve clinical representation learning. By fusing ECG signals with textual cardiac queries instead of fixed labels, SuPreME enables zero-shot classification of unseen conditions without further fine-tuning. We evaluate SuPreME on six downstream datasets covering 106 cardiac conditions, achieving superior zero-shot AUC performance of 77.20%, surpassing state-of-the-art eSSLs by 4.98%. Results demonstrate SuPreME’s effectiveness in leveraging structured, clinically relevant knowledge for high-quality ECG representations.
@@ 1 Introduction
To reduce dependence on annotations, recent advancements in ECG self-supervised learning (eSSL) have enabled the extraction of representative features from large-scale unannotated ECGs using contrastive or generative tasks (Eldele et al., 2021; Kiyasseh et al., 2021; Na et al., 2024). Despite their promise, these methods often rely on strong signal-level augmentations that may distort the semantic integrity of the signal and require complex pretext task designs (Kiyasseh et al., 2021). Multimodal learning approaches (Liu et al., 2024; Li et al., 2024) have also been proposed to learn ECG representations by leveraging free-text ECG reports. However, these methods face challenges due to noise in textual data and the complexities of language grammar, which can hinder learning efficiency (Wu et al., 2023).
Figure 1: Overview including (a) ECG clinical entity extraction with LLMs, (b) supervised pre-training with SuPreME, and (c) zero-shot prompted evaluation.
To address these limitations and develop a scalable, simple, and effective ECG pre-training framework, we propose SuPreME, a Supervised Pre-training framework for Multimodal ECG representation learning. Our contributions are threefold: (a) We introduce an automated pipeline that extracts high-quality clinical entities from raw ECG reports using an instruction-tuned LLM enriched with domain-specific knowledge (Figure 1(a)). Extracted entities are deduplicated and mapped to dataset-specific standardized diagnostic terms using clinician-validated resources (e.g., SNOMED CT, UMLS, SCP-ECG [^1] SNOMED CT and UMLS are standardized clinical terminology databases; SCP-ECG refers to the Standard Communication Protocol for Computer-Assisted Electrocardiography.), forming a dataset-specific global cardiac query list without manual annotation. This process enables scalable, consistent labeling and captures richer semantics than coarse-grained or free-text alternatives. (b) Leveraging these standardized cardiac queries, we propose SuPreME (Figure 1(b-c)), a multimodal framework that directly fuses ECG signals with cardiac queries through a lightweight Cardiac Fusion Network (CFN). Unlike prior methods (e.g., MERL) that rely on raw free-text inputs or handcrafted pretext tasks, SuPreME requires no signal-level augmentation or contrastive loss design, offering an efficient and interpretable multi-label supervision strategy grounded in standardized cardiac queries. (c) We pre-train SuPreME on 771,500 ECG signals paired with 295 global standardized cardiac queries from MIMIC-IV-ECG (Gow et al., 2023) (Appendix A.1.1). On six downstream datasets (e.g., PTB-XL, CPSC-2018, Chapman-Shaoxing-Ningbo; Appendix A.1.2), it achieves a new state-of-the-art zero-shot AUC of 77.20%, significantly outperforming existing eSSL and multimodal baselines, including those fine-tuned with 10–100% labeled data. SuPreME also shows strong data efficiency and generalization, with zero-shot performance under only 20% pre-training data surpassing fully fine-tuned eSSLs.
@@ 2 Related Work
ECG-Text Multimodal Learning. Multimodal learning has advanced significantly in biomedical applications, especially in vision-language pre-training (VLP) frameworks for radiology (Liu et al., 2023b, a; Wan et al., 2024; Zhang et al., 2023b; Wu et al., 2023; Abbaspourazad et al., 2023), which align radiology images with structured knowledge from reports to reduce noise and improve robustness. However, ECG-Text multimodal learning holds substantial potential for further development. Methods like MERL (Liu et al., 2024) and ECG-LM (Yang et al., ) integrate ECG signals and raw text reports but struggle with noise and inconsistencies in unstructured reports. Others, such as KED (Tian et al., 2024), use structured labels and contrastive learning strategies but face challenges from label noise and LLM-generated knowledge hallucinations. Our approach addresses these issues by structuring reports into meaningful entities, reducing noise, and aligning them with ECG signals without reliance on LLM-augmented content, minimizing hallucination risks while enabling efficient representation learning and downstream flexibility.
(a) Design of ECG report entity extraction with (i) knowledge-enhanced prompt engineering, and (ii) candidate entity deduplication and mapping.
(b) ECG 1D ViT encoder in the SuPreME, with both lead-wise and position embedding.
(c) Architecture of the Cardiac Fusion Network (CFN) in the SuPreME.
Figure 3: Implementation of the supervised ECG-Text multimodal pre-training framework including (a) ECG report entity extraction, (b) ECG 1D ViT encoder, and (c) architecture of the Cardiac Fusion Network.
@@ 3 Methodology
SuPreME extracts structured clinical entities from ECG reports via an instruction-tuned LLM (Section 3.1) to form cardiac queries, which are fused with ECG signals via Cardiac Fusion Network (CFN) in a shared latent space, enabling zero-shot classification of unseen cardiac conditions without fine-tuning (Section 3.3), thus yielding scalable, clinically meaningful representations.
3.1 LLM-based Clinical Entity Extraction
Enrich LLM with Domain Knowledge. ECG reports generated by 12-lead devices (Appendix A.2) contain diverse and nuanced descriptions of cardiac conditions. To contextualize these free-text reports for clinical entity extraction, we first construct a cardiac-specific vocabulary using GPT-4o-mini (Appendix A.3.1), by filtering, normalizing, and aggregating terminology from clinician-validated resources such as SNOMED CT, UMLS, and SCP-ECG (Bodenreider, 2004; Donnelly et al., 2006; Rubel et al., 2016). The vocabulary covers both complete diagnostic terms (e.g., sinus rhythm) and commonly used abbreviations (e.g., LVH for Left Ventricular Hypertrophy). This domain knowledge is then integrated into the LLM prompt design to guide the subsequent extraction process.
Knowledge-Guided Entity Extraction. We employ an instruction-tuned LLM to extract clinical entities from unstructured ECG reports. The model is prompted using structured instructions and few-shot examples (Appendix A.3.2), incorporating the curated cardiac vocabulary to enhance contextual understanding. It extracts diagnostic expressions (e.g., waveform patterns, cardiac abnormalities) along with their associated certainty. Extracted entities are categorized as Normal, Abnormal, or Uncertain, as illustrated in Figure 3(a). Entities marked with low certainty (e.g., containing “probably” or “cannot rule out”) are discarded to improve diagnostic precision. Within each report, semantically similar expressions are merged to prepare for cross-report alignment and standardization.
Entity Deduplication and Mapping. Although the extraction is guided and structured, lexical variability due to differing physician writing styles and device-specific formats still leads to redundant or inconsistent entities. To resolve this, we apply the curated cardiac vocabulary for entity standardization. Both extracted entities and reference terms are encoded using MedCPT, a medical BERT model pre-trained for clinical semantic similarity (Jin et al., 2023). Cosine similarity between embeddings is computed to perform soft matching. Entities that exhibit high average similarity to a reference term are aligned and deduplicated with a threshold selected and clinically validated by experienced cardiologists with over ten years of practice. Case study is provided in Appendix A.3.3.
@@ 3.2 Multimodal ECG Supervised Learning
ECG Embedding with Vision Transformer. The Vision Transformer (ViT) (Dosovitskiy et al., 2020), designed for 2D image processing, reshapes images into sequences of flattened patches for Transformer-based analysis. Similarly, ECG signals exhibit temporal and structural patterns analogous to the spatial relationships in images. We then adapt its architecture by dividing ECG time series into fixed-size patches, as shown in Figure 3(b).
@@
Multi-lead ECG signals are represented as $\mathbf{x}\in\mathbb{R}^{B\times L\times T}$, where $B$ is the batch size, $L$ the number of leads, and $T$ the number of time steps. Each lead’s signal is independently segmented into $N=T/P$ non-overlapping patches of length $P$, resulting in $\mathbf{x}_{i,j}\in\mathbb{R}^{B\times P}$ for lead $i$ and patch index $j$. Each patch is flattened and passed through a shared linear projection layer $\mathbf{W}_{p}\in\mathbb{R}^{P\times D}$ to produce a token embedding $\mathbf{z}_{i,j}\in\mathbb{R}^{B\times D}$:
$\begin{array}[]{c}\mathbf{z}_{i,j}=\mathbf{x}_{i,j}\mathbf{W}_{p}\\[3.00003pt] \mathbf{z^{\prime}}_{i,j}=\mathbf{z}_{i,j}+\mathbf{e}_{i}+\mathbf{p}_{j}\end{array}$ (1)
To preserve lead-specific features and spatial-temporal information, we introduce a unique learnable lead embedding $\mathbf{e}_{i}\in\mathbb{R}^{D}$ for each lead $i$, and a positional embedding $\mathbf{p}_{j}\in\mathbb{R}^{D}$ for each patch index $j$. These embeddings are element-wise added to the projected patch token $\mathbf{z}_{i,j}$, forming the enriched representation $\mathbf{z^{\prime}}_{i,j}$. All enriched tokens from all leads and patches are then concatenated to form the Transformer encoder input sequence:
$\begin{array}[]{c}\mathbf{Z}=[\mathbf{z^{\prime}}_{1,1},\dots,\mathbf{z^{\prime}}_{1,N},\dots,\mathbf{z^{\prime}}_{L,N}]\\[3.00003pt] \mathbf{F}_{\text{ECG}}=\text{MLP}_{\text{ECG}}(\text{Dropout}(\mathbf{Z}))\end{array}$ (2)
The final token sequence $\mathbf{Z}\in\mathbb{R}^{B\times(L\cdot N)\times D}$ is passed through Transformer encoders to extract high-level ECG representations. Each block consists of multi-head self-attention and feed-forward sublayers, with residual connections and layer normalization. To enhance generalization, we apply stochastic depth dropout to the residual paths.
Before multimodal fusion, the output features are passed through a modality-specific two-layer multilayer perceptron (MLP) projection head with an intermediate non-linear activation. This projection maps the ViT output from its internal width $D$ to a shared latent dimension $D^{\prime}$ aligned with the textual modality, forming representation $\mathbf{F}_{\text{ECG}}\in\mathbb{R}^{B\times(L\cdot N)\times D^{\prime}}$ used as the CFN input. Implementation details are provided in Appendix A.4.1,  A.4.2.
Cardiac Query Embedding with MedCPT. Instead of relying on fixed categorical labels, our framework adopts a flexible and semantically meaningful approach based on textual cardiac queries. For the pre-training dataset, we construct fine-grained diagnostic queries by applying the LLM-based entity extraction pipeline described in Section 3.1. Specifically, we generate a dataset-specific global query list of size $M$, consisting of standardized cardiac terms derived from the deduplicated and mapped output of the extraction process.
@@
$\begin{array}[]{c}\mathbf{E}[i,:]=\text{Trm}([\text{CLS}]\ q_{i}\ [\text{SEP}])\\[3.00003pt] \mathbf{F}_{\text{Query}}=\text{MLP}_{\text{Query}}(\text{Dropout}(\mathbf{E}))\end{array}$ (3)
All $M$ query embeddings $\mathbf{E}\in\mathbb{R}^{M\times 768}$ are then passed through a modality-specific two-layer MLP projection head with an intermediate activation function to obtain the final representations $\mathbf{F}_{\text{Query}}\in\mathbb{R}^{M\times D^{\prime}}$ in the shared $D^{\prime}$-dimensional latent space aligned with ECG token representations. Further implementations are provided in Appendix A.4.3.
Alignment by Cardiac Fusion Network. The Cardiac Fusion Network (CFN) fuses ECG signals with textual cardiac queries using a multi-layer Transformer decoder architecture, where query embeddings act as decoder inputs and ECG features serve as the encoder memory, following a standard cross-attention formulation (Figure 3(c)).
Given a batch of ECG features $\mathbf{F}_{\text{ECG}}\in\mathbb{R}^{B\times(L\cdot N)\times D^{\prime}}$ and query embeddings $\mathbf{F}_{\text{Query}}\in\mathbb{R}^{M\times D^{\prime}}$, CFN fuses the two modalities through cross-attention. During pre-training, each ECG sample is paired with the same $M$ cardiac query embeddings, allowing CFN to learn a joint representation that captures query-conditioned patterns in the signal. The decoder attends to ECG patterns while grounding the prediction in the semantics of each diagnostic query, outputting $\mathbf{H}\in\mathbb{R}^{B\times M\times D^{\prime}}$ which is passed through a single MLP classification head shared across all queries, producing $M$ binary logits per ECG sample, where each logit indicates the relevance of a specific query to the signal input:
$\mathbf{Logits}=\text{MLP}_{\text{CFN}}(\mathbf{H})\in\mathbb{R}^{B\times M}$ (4)
In pre-training, we supervise CFN using weak binary labels derived from the entity extraction pipeline. Each ECG report is matched against the global query list of $M$ standardized diagnostic terms. A binary label of 1 is assigned if a mapped report entity matches a query; otherwise 0. This results in a sparse $M$-dimensional multi-label vector per ECG sample. To avoid data leakage and ensure modality separation, raw ECG reports are never used directly as input to the model. Instead, diagnostic query list serves as input with queries embedded independently and applied uniformly to every ECG sample.
This formulation enables SuPreME to perform open-set classification with a flexible, scalable query interface, supporting multi-label learning while maintaining clear supervision-query decoupling. CFN initialization is in Appendix A.4.4.